# Adapting Noise to Data: Generative Flows from 1D Processes

### v3

@@ 2.3 Stochastic Processes and Velocity Fields
Then $f(t)\coloneqq 1-t$, $g(t)\coloneqq t^{2}$ yields (independent) FM [lipman2023flow], and $f(t)\coloneqq\exp\left(-\frac{h(t)}{2}\right)$, $g(t)\coloneqq 1-\exp\left(-h(t)\right)$, where $h(t)\coloneqq\int_{0}^{t}\beta_{\rm min}+s(\beta_{\rm max}-\beta_{\rm min})\mathop{}\!\mathrm{d}s$ with, e.g., $\beta_{\rm min}=0.1,\,\beta_{\rm max}=20$, corresponds to processes used in score-based generative models [song2021scorebasedgenerativemodelingstochastic], see Appendix B.
Remark 2 (Optimal Coupling).
Instead of considering (possibly independent) random variables $\mathbf{X}_{0},\mathbf{X}_{1}$ and their induced processes such as in (8), we can also introduce a coupling between them. Let $X_{0}\sim\mu_{0},X_{1}\sim\mu_{1}$ and consider the optimal coupling $\pi\in\Pi_{o}(\mu_{0},\mu_{1})$. Then define the induced curve $(e_{t})_{\sharp}\pi$, where $e_{t}(x,y):=(1-t)x+ty$. This curve is a geodesic between $\mu_{0}$ and $\mu_{1}$ in the Wasserstein geometry, for more details see [WS2025]. Using the optimal coupling $\pi\in\Pi_{o}(\mu_{0},\mu_{1})$ yields an Optimal Transport FM objective
$\mathcal{L}_{\mathsf{OT\text{-}CFM}}(\theta)=\mathbb{E}_{t\sim\mathcal{U}(0,1),\,(x,y)\sim\pi}\Big[\big\|v_{\theta}\big((1-t)x+ty,\,t\big)-(y-x)\big\|_{2}^{2}\Big].$ (10)
The velocity field minimizing this objective then satisfies the continuity equation together with $e_{t}\sharp\pi$. In contrast to using the independent coupling, this can lead to reduced variance in training and both shorter and straighter paths, see [tong2024improvinggeneralizingflowbasedgenerative, pooladian2023multisampleflowmatchingstraightening].
Motivated by the fact that a multi-dimensional Wiener process $\mathbf{W}_{t}\in\mathbb{R}^{d}$ consists of independent (and identically distributed) 1D components $\mathbf{W}_{t}=(W_{t}^{1},...,W_{t}^{d})$, we propose to construct a $d$-dimensional flow $\mathbf{Y}_{t}$ componentwise, based on independent one-dimensional processes $Y_{t}^{i}$.
3 From One-Dimensional to Multi-Dimensional Flows
Restricting ourselves to processes $\mathbf{Y}_{t}$ that decompose into one-dimensional components allows us to propose a general construction method for accessible conditional flows in FM. Let $Y_{t}^{1},\dots,Y_{t}^{d}$ be a family of independent one-dimensional stochastic processes with time dependent laws $\mu_{t}^{i}\in\mathcal{P}_{2}(\mathbb{R})$. For each $i=1,\dots,d$, let $v_{t}^{i}\colon\mathbb{R}\to\mathbb{R}$ be the associated velocity field such that the pair $(\mu_{t}^{i},v_{t}^{i})$ satisfies the one-dimensional continuity equation (1). Define the product measure $\mu_{t}\in\mathcal{P}_{2}(\mathbb{R}^{d})$ by
$\mu_{t}(x)\;=\;\prod_{i=1}^{d}\mu_{t}^{i}(x^{i}),\qquad x=(x^{1},\dots,x^{d})\in\mathbb{R}^{d}.$ (11)
For the $d$-dimensional process $\mathbf{Y}_{t}\coloneqq(Y_{t}^{1},\dots,Y_{t}^{d})$, independence implies that its law is exactly $\mu_{t}$. Moreover, by the following proposition, the corresponding $d$-dimensional velocity field is given componentwise, see [DCFS2025].
Proposition 3.
Let $\mu_{t}$ be given by (11), where the $\mu^{i}_{t}$ are absolutely continuous curves in $\mathbb{R}$ with velocity fields $v^{i}_{t}$. Then $\mu_{t}$ satisfies a multi-dimensional continuity equation (1) with a velocity field which decomposes into the univariate velocities
$v_{t}(x):=\left(v^{1}_{t}(x^{1}),\dots,v^{d}_{t}(x^{d})\right).$ (12)
Therefore, as long as we have access to the velocity field associated to our one-dimensional processes, we can construct accessible conditional flows for FM:
@@ 3 From One-Dimensional to Multi-Dimensional Flows
    One-dimensional noise: Start with a one-dimensional process and an associated absolutely continuous curve $\mu_{t}$ with $\mu_{0}=\delta_{0}$, $0\in\mathbb{R}$, where you can compute the velocity field $v_{t}$ in the 1D continuity equation
@@
    Multi-dimensional noise: Set up a multi-dimensional conditional flow model starting in $\mu_{0}=\delta_{0}$, $0\in\mathbb{R}^{d}$ with possibly different, but independent 1D processes as described in Section 3.
@@
To outline the use of this recipe, we explore three interesting 1D (noising) processes $Y_{t}$ in connection with their respective PDEs, for which our approach via reduction to one dimension is nicely applicable, namely the
@@
  • Uniform process $U_{t}$ and the gradient flow of the maximum mean functional $\mathcal{F}_{\nu}\coloneqq{\rm MMD}_{K}(\cdot,\nu)$ with negative distance kernel K(x,y)=|xy|K(x,y)=-|x-y| and $\nu=\mathcal{U}(-b,b)$.
Paths of the processes are depicted in Figure 2.
@@
In each case, the absolutely continuous curve starting in $\delta_{0}$ and the corresponding velocity field can be calculated analytically. Note that in contrast to the Wiener process $W_{t}$ usually seen in diffusion and flow matching models, the latter two processes $K_{t},U_{t}$ do not enjoy a trivial analogue in multiple dimensions: in case of $K_{t}$ the corresponding PDE (damped wave equation) is no longer mass-conserving in dimension $d\geq 3$, see [TL2016]; in case of $U_{t}$ the mere existence of the MMD gradient flow in multiple dimensions is unclear by the lack of convexity of the MMD, see [HGBS2022]. Our general construction method makes these 1D processes accessible for generative modeling in arbitrary dimensions.

### v4

@@ 2.4 Velocity Fields via Stochastic Processes
Then $f(t)\coloneqq 1-t$, $g(t)\coloneqq t^{2}$ yields (independent) FM (Lipman et al., 2023), and $f(t)\coloneqq\exp\left(-\frac{h(t)}{2}\right)$, $g(t)\coloneqq 1-\exp\left(-h(t)\right)$, where $h(t)\coloneqq\int_{0}^{t}\beta_{\rm min}+s(\beta_{\rm max}-\beta_{\rm min})\,\mathrm{d}s$ with, e.g., $\beta_{\rm min}=0.1,\,\beta_{\rm max}=20$, corresponds to processes used in score-based generative models (Song et al., 2021), see Appendix B.
3 Reduction to One-Dimensional Flows
Recall that in standard FM and diffusion models, isotropic Gaussian noise is gradually added to the data, i.e.
$\mathbf{X}_{t}=\alpha_{t}\mathbf{X}_{0}+\mathbf{N}_{t},$ (12)
where $\mathbf{N}_{t}=(N_{t}^{1},\dots,N_{t}^{d})$ with i.i.d. $N_{t}^{i}\sim\mathcal{N}(0,\sigma_{t}^{2})$. Motivated by this, we restrict ourselves to noising processes $\mathbf{N}_{t}$ that decompose into one-dimensional components. This allows us to propose a general construction method for accessible conditional flows in FM.
Let $N_{t}^{1},\dots,N_{t}^{d}$ be a family of independent 1D stochastic processes with laws $\mu_{t}^{i}\in\mathcal{P}_{2}(\mathbb{R})$. For each $i=1,\dots,d$, let $v_{t}^{i}\colon\mathbb{R}\to\mathbb{R}$ be the associated velocity field such that the pair $(\mu_{t}^{i},v_{t}^{i})$ satisfies the one-dimensional continuity equation (2). Then, the law of the $d$-dimensional process $\mathbf{N}_{t}\coloneqq(N_{t}^{1},\dots,N_{t}^{d})$ is the product measure
$\mu_{t}(x)\;=\;\prod_{i=1}^{d}\mu_{t}^{i}(x^{i}),\quad x=(x^{1},\dots,x^{d})\in\mathbb{R}^{d},$ (13)
and the corresponding $d$-dimensional velocity field is given componentwise, see (Duong et al., 2025).
Proposition 3.1.
Let $\mu_{t}$ be given by (13), where the $\mu^{i}_{t}$ are absolutely continuous curves in $\mathbb{R}$ with velocity fields $v^{i}_{t}$. Then $\mu_{t}$ satisfies a multi-dimensional continuity equation (2) with a velocity field which decomposes into the univariate velocities $v_{t}(x):=\left(v^{1}_{t}(x^{1}),\dots,v^{d}_{t}(x^{d})\right)$.
In summary, assuming access to the 1D velocities, we can construct accessible conditional flows for FM as follows:
@@ 3 Reduction to One-Dimensional Flows
    One-dimensional noise: Start with a 1D process and an associated curve $\mu_{t}$ with $\mu_{0}=\delta_{0}$, $0\in\mathbb{R}$ and an accessible velocity field $v_{t}$ in the 1D continuity equation
@@
    Multi-dimensional noise: Set up a multi-dimensional conditional flow model starting in $\mu_{0}=\delta_{0}$, $0\in\mathbb{R}^{d}$ with possibly different, but independent 1D processes as described in (13).
@@
Remark 3.2 (Learning Cross Dimensional Correlation of the Data.).
As usual in diffusion or flow matching models, the velocity field (6) (or score function in diffusion) is responsible for learning cross-dimensional correlations and dependencies of the target $\mathbf{X}_{0}$.
Concerning step 1, we explore three interesting 1D (noising) processes in connection with their respective PDEs in the Appendix A, namely:
@@
- Uniform process $U_{t}$ and MMD gradient flow.
@@
In each case, we explicitly calculate the respective conditional measure flow and its conditional velocity field in Appendix A, such that the conditional flow matching loss $\mathcal{L}_{\text{CFM}}$ can be minimized. In each case, the absolutely continuous curve starting in $\delta_{0}$ and the corresponding velocity field can be calculated analytically. In contrast to usual the Wiener process $W_{t}$ in diffusion and FM models, the latter two processes $K_{t},U_{t}$ do not enjoy a trivial analogue in multiple dimensions: in case of $K_{t}$ the corresponding PDE (damped wave equation) is no longer mass-conserving in dimension $d\geq 3$, see (Tautz and Lerche, 2016); in case of $U_{t}$ the mere existence of the MMD gradient flow in multiple dimensions is unclear by the lack of convexity of the MMD, see (Hertrich et al., 2024). Our general construction method makes these 1D processes accessible for generative modeling in arbitrary dimensions, hinting at a wide range of suitable noising processes. For example, the recent work (Han et al., 2025) utilizes the Kac framework for model distillation.
As an alternative to above componentwise construction, we describe an approach using radially symmetric processes in Appendix D.
Figure 3: A generated trajectory from the learned quantile latent (left) to the unevenly weighted Gaussian mixture target (right). The adapted latent $\mathbf{Q}_{\phi}$ (left image) is already close to the target distribution $\mathbf{X}_{0}\sim\mu_{0}$.
4 One-Dimensional Flows with Quantiles
We have outlined how to construct accessible conditional flows for FM using 1D noising processes. So far, we have considered three fixed 1D processes. However, every 1D distribution can be described by its quantile functions. This will enable us in Section 5 to construct arbitrary noising processes, and to learn the noise.
To this end, we revisit the connection between 1D distributions and their quantile functions. Furthermore, we introduce quantile processes and quantile interpolants.
4.1 Background on Quantile Functions
The restriction to componentwise noising processes $\mathbf{N}_{t}=(N_{t}^{1},\dots,N_{t}^{d})$ in (12) [^3] Besides componentwise 1D processes we may also use triangular decompositions, not addressed in this paper. allows us to use the quantile functions of the 1D components. The cumulative distribution function (CDF) $R_{\mu}$ of $\mu\in\mathcal{P}_{2}(\mathbb{R})$ and its quantile function $Q_{\mu}$ are given by
$R_{\mu}(x)$ $\coloneqq\mu\big((-\infty,x]\big),\quad x\in\mathbb{R},$ (15)
$Q_{\mu}(u)$ $\coloneqq\min\{x\in\mathbb{R}:R_{\mu}(x)\geq u\},\quad u\in(0,1).$ (16)
The quantile functions form a closed, convex cone
$\mathcal{C}\coloneqq\{f\in L_{2}(0,1):f\text{ increasing }a.e.\}$ (17)
in $L_{2}(0,1)$. The mapping $\mu\mapsto Q_{\mu}$ is an isometric embedding of $(\mathcal{P}_{2}(\mathbb{R}),W_{2})$ into $(L_{2}(0,1),\|\cdot\|_{L_{2}})$, meaning that
$W_{2}^{2}(\mu,\nu)=\int_{0}^{1}\big|Q_{\mu}(s)-Q_{\nu}(s)\big|^{2}\,\,\mathrm{d}s$ (18)
and $\mu=Q_{\mu,\sharp}\mathcal{U}_{(0,1)}$. Let $U\sim\mathcal{U}[0,1]$ be uniformly distributed on $[0,1]$. Now, any probability measure flow $\mu_{t}$ can be described by their quantile flow $Q_{t}\coloneqq Q_{\mu_{t}}$, such that $\mu_{t}=Q_{t,\sharp}\mathcal{U}_{(0,1)}$ and $Q_{t}\circ U$ is a stochastic process with marginals $\mu_{t}$.
Parameterization. Quantile functions are both universal by being able to express any 1D distribution, and simple to parametrize, since they only need to satisfy the requirements from (17). For example quantile functions can be modeled using rational quadratic splines (Gregory and Delbourgo, 1982; Durkan et al., 2019), as described in Section 6.1. Moreover, Dirac measures $\delta_{a}$ are easily represented in this framework via constant quantile functions $Q\equiv a$.
4.2 Quantile Processes
We can therefore model any multi-dimensional noising process, that decomposes into its components $\mathbf{N}_{t}=(N_{t}^{1},\dots,N_{t}^{d})$, via quantile functions. Namely let $X_{0}$ be any component $\mathbf{X}_{0}^{i}$ of $\mathbf{X}_{0}\sim\mu_{0}$, and $f,g:[0,1]\to\mathbb{R}$ smooth schedules fulfilling (8). We assume that we are given a flow $(Q_{t})_{t}$ of quantile functions $Q_{t}:(0,1)\to\mathbb{R}$, $t\in[0,1]$, which fulfill $Q_{0}\equiv 0$. Let $U\sim\mathcal{U}(0,1)$. We introduce the quantile process
$Z_{t}=f(t)X_{0}+Q_{g(t)}(U),\quad t\in[0,1].$ (19)
By choosing the quantile process $Q_{t}^{i}(U)\overset{d}{=}t\cdot Q_{\mathcal{N}(0,1)}(U)$, the resulting process coincides with the standard Gaussian interpolation process in (12).
4.3 Quantile Interpolants.
Assuming we are invertible on their respective image with the inverse given by the CDF $R_{t}:Q_{t}(0,1)\to\mathbb{R}$. Let us briefly mention how our setting fits into the framework of consistency models. To this end, we define the quantile interpolants for $s,t\in[0,1]$ by
$I_{s,t}(x,y)=f(s)x+Q_{g(s)}\big(R_{g(t)}(y-f(t)x)\big).$ (20)
This generalizes the interpolants used in Denoising Diffusion Implicit Models (DDIM), see Remark C.1.
Proposition 4.1.
$I_{s,r}(x,I_{r,t}(x,y))=I_{s,t}(x,y).$ (21)
The proof is given Appendix C. Proposition 4.1 allows us to also apply the concept of consistency models to our quantile process (19). In the Appendix C, we demonstrate this by means of the recently proposed inductive moment matching (IMM) (Zhou et al., 2025).
5 Adapting Noise to Data
As noted in many papers, see also the related work in Section 7, the choice of the noise distribution can have a significant impact on sampling performance. In Figure 1, we visualize the path length under the learned vector field for the checkerboard distribution. In Figure 2, we evaluate the Neals funnel (Neal, 2003) distribution under different noise distributions, ranging from compact-support to heavy-tailed.
Rather than manually selecting a noise family and tuning its parameters, we propose to learn the noise directly from the data. We focus on the simple setting $\mathbf{Q}_{g(t)}=g(t)\,\mathbf{Q}$ and schedules f(t)=1tf(t)=1-t, $g(t)=t$. This corresponds to the standard linear interpolation in FM with conditional velocity $v^{\mathbf{Q}}_{t}(x)=x/t$; see Appendix A.4.
We restrict to noise $\mathbf{Q}$ with independent components having law from the set
$S\coloneqq\big\{\nu\in\mathcal{P}_{2}(\mathbb{R}^{d}):\nu(x)=\prod_{i=1}^{d}\nu^{i}(x^{i})\big\},$
corresponding to quantile processes of the form
$X_{t}^{i}=(1-t)\,X_{0}^{i}+t\,Q^{i}(U^{i}),\quad t\in[0,1],$ (22)
with $\mathbf{Q}(\mathbf{u}):=(Q^{1}(u^{1}),\ldots,Q^{d}(u^{d}))$. The quantile functions $Q^{i}$ determine the scale, support, and tail behavior of each marginal of $\mathbf{Q}(\mathbf{U})$.
Learning Objective. We learn the noise $\mathbf{Q}_{\phi}$ by minimizing the Wasserstein distance between $\mu_{0}$ and $\nu_{\phi}:=(\mathbf{Q}_{\phi})_{\#}\mathcal{U}([0,1]^{d})$,
$\mathcal{L}_{\mathrm{AN}}(\phi)=W_{2}^{2}(\mu_{0},\nu_{\phi}).$ (23)
Note that due to the restriction of our quantiles to the class $S$, the minimizer of (23) is in general not $\mu_{0}$. Crucially, the independence constraint restricts $(\mathbf{Q}_{\phi})_{\#}\,\operatorname{\mathcal{U}}([0,1]^{d})$ to per-coordinate adaptation and prevents encoding cross-dimensional correlations. The latter are introduced via the optimal transport coupling $(x,y)$ and modeled by the velocity field through the target (yx)(y-x). By this separation the latent remains simple and computationally efficient, while delegating dependencies to the flow.
Joint Training Objective. Although our quantiles can be trained independently, in order to provide an aligned training signal for the velocity field, we propose to train $\mathbf{Q}_{\phi}$jointly with velocity $v^{\theta}$. Hence, we aim to minimize the loss
$\mathcal{L}(\theta,\phi)$ $=\mathcal{\mathcal{L}}_{\mathsf{CFM}}(\theta,\phi)+\lambda\,\mathcal{\mathcal{L}}_{\operatorname{AN}}(\phi),\quad\lambda>0,$ (24)
with ${\mathcal{L}}_{\mathsf{CFM}}(\theta,\phi)$ given by
$\mathbb{E}_{t,(x,y_{\phi})\sim\pi_{\phi}}\Big[\big\|v^{\theta}_{t}\big((1-t)x+ty_{\phi}\big)-(\operatorname{{sg}}(y_{\phi})-x)\big\|_{2}^{2}\Big],$ (25)
where $\pi_{\phi}\in\Pi_{o}(\mu_{0},\nu_{\phi})$ and $\nu_{\phi}$ and $\operatorname{sg}(\cdot)$ denotes the stop-gradient operator. We visualize the effect of choosing $\lambda=0$ in Figure 18.
Implementation Details. In practice, we optimize via minibatches; see Appendix F.2 for details. We compute a minibatch OT map $T$ minimizing $\sum_{j=1}^{B}\|\mathbf{x}_{0}^{(j)}-\mathbf{y}^{(T(j))}\|_{2}^{2}$ for batches from $\mathbf{X}_{0}$ and $\mathbf{Q}_{\phi}(\mathbf{U})$. Crucially, this coupling is reused both for optimizing the quantile and for OT-FM, see Remark 2.1. We train the quantile jointly for a fixed number of iterations before freezing its parameters; thereafter only the velocity field is optimized.
Algorithm 1 Joint learning of 1D quantiles and FM velocity
Input: dataset $\mathcal{D}$, batch size $B$, weight $\lambda$, iterations $K$; quantile model $\mathbf{Q}_{\phi}$, velocity model $v^{\theta}$
for $k=1$ to $K$ do
  Sample $\{\mathbf{x}_{i}\}_{i=1}^{B}\sim\mathcal{D}$, $\{\mathbf{u}_{j}\}_{j=1}^{B}\sim\mathcal{U}([0,1]^{d})$, $\{t_{j}\}_{j=1}^{B}\sim\mathcal{U}(0,1)$
  $C_{ij}\leftarrow\|\mathbf{x}_{i}-\mathbf{Q}_{\phi}(\mathbf{u}_{j})\|_{2}^{2}$
  $T\leftarrow\arg\min_{T}\sum_{i=1}^{B}C_{i,T(i)}$
  Define $P$ such that $P(j)=i$ iff $T(i)=j$
  for $j=1$ to $B$ do
   $\mathbf{z}_{j}\leftarrow(1-t_{j})\mathbf{x}_{P(j)}+t_{j}\,\mathbf{Q}_{\phi}(\mathbf{u}_{j})$
   $\mathbf{v}_{\mathrm{target},j}\leftarrow\operatorname{sg}\!\big(\mathbf{Q}_{\phi}(\mathbf{u}_{j})-\mathbf{x}_{P(j)}\big)$
  end for
  $\widehat{\mathcal{L}}_{\mathrm{AN}}\leftarrow\frac{1}{B}\sum_{j=1}^{B}\|\mathbf{x}_{P(j)}-\mathbf{Q}_{\phi}(\mathbf{u}_{j})\|_{2}^{2}$
  $\widehat{\mathcal{L}}_{\mathsf{CFM}}\leftarrow\frac{1}{B}\sum_{j=1}^{B}\|v^{\theta}(\mathbf{z}_{j},t_{j})-\mathbf{v}_{\mathrm{target},j}\|_{2}^{2}$
  $\widehat{\mathcal{L}}\leftarrow\widehat{\mathcal{L}}_{\mathsf{CFM}}+\lambda\,\widehat{\mathcal{L}}_{\mathrm{AN}}$
  Update $(\theta,\phi)$ by a gradient step on $\widehat{\mathcal{L}}$
end for
Output: learned parameters $(\theta,\phi)$
6 Experiments
To provide intuition and validate our proposed method, we conduct experiments on both synthetic and image datasets. First we briefly outline how to efficiently parametrize quantile functions.
6.1 Rational Quadratic Splines
We parameterize each quantile function $Q^{i}$ using a rational quadratic spline (RQS) (Gregory and Delbourgo, 1982; Durkan et al., 2019). Unlike normalizing flows that compose many RQS layers within coupling architectures, we use a single RQS per coordinate with directly optimized parameters.
For coordinate $i$, the spline $S^{i}_{\phi}:[-B,B]\to[-B,B]$ is defined by $K$ knots with associated bin widths $\{w_{k}\}_{k=1}^{K}$, bin heights $\{h_{k}\}_{k=1}^{K}$, and slopes $\{s_{k}\}_{k=0}^{K}$. We enforce positivity via softplus and normalize widths and heights to span $2B$. This guarantees strict monotonicity. Outside [B,B][-B,B], we extend $S^{i}_{\phi}$ linearly with $C^{1}$ continuity at the boundaries.
To map from $(0,1)$ to the spline domain, we apply $\psi(u)=\operatorname{logit}(u)$ or $\psi(u)=B(2u-1)$, yielding the quantile
$Q^{i}_{\phi}(u)=a_{i}\cdot S^{i}_{\phi}(\psi(u))+b_{i},$ (26)
where the scale $a_{i}>0$ and bias $b_{i}$ are learned per coordinate.
The total parameter count is $\mathcal{O}(Kd)$: for $K=32$ bins and $d=3072$ (CIFAR-10), approximately 300k parameters. This lightweight parametrization, combined with reusing the minibatch OT coupling and freezing the quantile after 55k iterations, introduces minimal overhead compared to standard Gaussian OT-FM. On CIFAR-10 (NVIDIA RTX 5090), we measure approximately $2.7\%$ overhead during joint training and $0.5\%$ after freezing.
6.2 Synthetic Datasets
We begin by qualitatively analyzing our algorithm on several synthetic 2D distributions, see also Appendix F.1, each designed to highlight a specific aspect of our approach. We provide intuition about the learned latent distribution and demonstrate that it is closer to the data in the Wasserstein sense, yields shorter transport paths, and successfully captures the tail behavior.
Gaussian Mixture Model (GMM).
We first consider a 2D GMM with nine unevenly weighted modes, as visualized in Figure 3. Due to the independence assumption inherent in our factorized quantile function, the learned latent cannot perfectly replicate the target’s joint distribution and is not the product of the correct marginals, see also Example  E.1. Instead, it approximates a distribution where the components cannot further independently improve the transport cost to the target.
Funnel Distribution.
The funnel distribution (Figure 2) presents a challenge due to its heavy-tailed, conditional structure. Several methods have been proposed in the diffusion context (Pandey et al., 2024; Shariatian et al., 2025b, a). We compare to (Pandey et al., 2024), where Student-$t$ parameters were hand-selected per dimension. Using a capacity-constrained network (three layers, width 64, no positional embeddings), we observe that a compact latent performs worst, followed by the Gaussian, while our learned latent successfully adapts to the target’s heavy tails (see Figure 12). [^4] Due to the high variance when sampling minibatch from the funnel, we pre-train the quantile and use the independent coupling for all models.
Checkerboard Distribution.
The checkerboard in Figure 11 features compact support. Our method learns a latent approximating a uniform distribution over the target’s support. Combined with OT coupling, the resulting transport paths are substantially shorter than starting from a Gaussian as visualized in Figure 1. In addition, in training, the vector field converges faster, see Figure 13. This underscores our central claim: combining a data-dependent latent with a data-dependent coupling can significantly improve performance.
6.3 Image Datasets
Next, we analyze our method on standard image generation benchmarks. Our quantile is extremely lightweight compared to the U-Net architecture used for the flow model. In high-dimensional settings and given fixed batch sizes, the signal for the quantile function can be noisy, potentially leading to degenerate solutions. To mitigate this, we add a regularization term to the loss that penalizes the expected negative log-determinant of the Jacobian of the quantile. Since the quantile maps from a uniform distribution $U\sim\mathcal{U}[0,1]$, this term equals the negative differential entropy $h$ of the learned latent
$h(Q(U))$ $=h(U)+\mathbb{E}[\log|\det J_{Q}(U)|]$
$=\mathbb{E}[\log|\det J_{Q}(U)|].$
Access to analytic derivatives makes this efficient, for more details see E.2. In total, this yields the loss $\mathcal{L}(\theta,\phi)=\mathcal{\mathcal{L}}_{\mathsf{CFM}}(\theta,\phi)+\lambda{\mathcal{L}}_{\operatorname{AN}}(\phi)+\beta\mathbb{E}[\log|\det J_{Q}(U)|]$.
Figure 4: Ablation on U-Net capacity for MNIST using channels $8,16,32$. The FID curves show that our method achieves significantly lower FIDs when using less channels.
Figure 5: Left to Right: Generated samples, samples from the learned latent and mean and standard deviation of the learned latent.
Baseline
$\beta=0.8$
Euler steps $\rightarrow$ 20 100
$\boldsymbol{\beta}$-logdet $\downarrow$ FID FID
0.2 7.81 4.75
0.3 7.48 4.53
0.5 7.66 4.49
0.8 7.77 4.42
1.0 8.35 4.66
Baseline 8.42 4.63
Figure 6: CIFAR results for a selection of regularization parameters and for the baseline, for complete results see Figure 17. Our method reached the best validation FID after 320k steps, while the baseline took 340k. We evaluated the FID using 5 seeds and report the mean. We used those checkpoints for the evaluation. The visualized samples were generated using 100 Euler steps.
MNIST.
MNIST exhibits strong marginal structure: center pixels are frequently active while border pixels are nearly always zero. Our learned quantile captures these statistics, concentrating mass in regions corresponding to active pixels as shown in Figure 5. Where a pixel is essentially always black, the learned quantile concentrates around that value; in center regions with higher uncertainty, the quantiles remain spread around zero (gray), accurately reflecting the data variability. We compare learned and empirical quantiles at different pixel locations in Figure 15.
This dataset is well-suited for our method since the marginal structure provides a strong learning signal. In Figure 4, we compare performance under capacity constraints using channels $8,16,32$ ($\lambda=1$, $\beta=0.1$). Our learned latent achieves significantly lower FIDs across all capacity levels. See Figure 16 in the Appendix for the total amount of parameters used for each model. By removing redundant information in the noise distribution, the network can allocate its parameters more efficiently toward modeling the relevant structure. The independence assumption prevents capturing spatial correlations (e.g., digit shapes).
CIFAR-10.
We evaluate on CIFAR-10 to demonstrate that our method scales to natural images and integrates seamlessly with standard architectures. We follow the setup of (Tong et al., 2024), using the U-Net architecture from (Nichol and Dhariwal, 2021), to enable a fair comparison. Figure 6 reports results for varying $\beta$ at fixed $\lambda=1$. Our method successfully learns a data-adapted latent while maintaining competitive FID scores, confirming that the quantile training remains stable at scale. See also Figure 18 for details on training stability. Improvements are marginal as expected, since CIFAR-10 features strong spatial and inter-channel correlations that a product measure cannot fully capture.
7 Related Work
Generative Models.
Current state-of-the-art methods largely rely on corrupting images with Gaussian noise. Diffusion models (Song et al., 2021; Ho et al., 2020; Song et al., 2022) and the closely related flow matching models (Liu et al., 2023; Lipman et al., 2023; Albergo et al., 2023) form the basis of most approaches. To further improve inference speed, few-step methods have achieved remarkable results (Song et al., 2023; Geng et al., 2025; Zhou et al., 2025).
Learning the Flow.
The following papers keep the Gaussian distribution as a starting point, but updated the interpolation path. There exist only few approaches to learn the noising process, (Bartosh et al., 2025) fit the forward diffusion process to the backward via a learned invertible map that is trained end-to-end, (Kapusniak et al., 2024) use metric flow matching, i.e., a neural network to adapt the path to a underlying Riemannian metric. In a related approach (Sahoo et al., 2024) learns a input-conditioned componentwise Gaussian noise schedule. In the setting of sampling from unnormalized target densities, (Blessing et al., 2025a) learn the latent noise by optimizing the mean and covariance of a Gaussian prior. A Gaussian mixture prior has been used both for sampling (Blessing et al., 2025b) and for generative modeling (Issachar et al., 2025).
Along the lines of normalizing flows and VAEs, there is a large body of work on learning more flexible noise distributions and priors (Stimper et al., 2022; Bauer and Mnih, 2019; Hickling and Prangle, 2025; Amiri et al., 2024).
Limitations of Gaussian Noise.
From a complementary perspective, (Wiese et al., 2019) propose to separate marginal modeling from dependence structure using copula and marginal flows, recognizing that standard architectures struggle with tail asymptotics, a motivation conceptually aligned with our componentwise quantile approach. On the other hand (Pandey et al., 2024; Zhang et al., 2024) design heavy-tailed diffusions using Student-$t$ latent distributions, and (Shariatian et al., 2025a) extend the framework to the family of $\alpha$-stable distributions.
Recent work by (Ghane et al., 2025) explains why heavy-tailed sampling can be challenging: diffusion models driven by Gaussian noise satisfy a concentration-of-measure property. In particular, when initialized from a Gaussian distribution, the generated distribution inherits many Gaussian-like properties. Moreover, by (Tam and Dunson, 2025), GANs, VAEs and diffusion models with Gaussian or log-concave noise distributions can only generate light-tailed samples and are not universal generators. See Figure 2 for a heavy-tailed example, where Gaussian noise fails.
8 Conclusions
We provide a ”quantile sandbox” for building generative models: a unifying theory and a practical toolkit that turns noise selection into a data-driven design element. Our construction plugs seamlessly into standard objectives including flow matching and few-step models. Furthermore, our experiments demonstrate that it is possible to learn extremely efficient, freely parametrized latent distributions beyond the usual smooth transformations of Gaussians. Our work opens promising directions for future research. Extensions include developing well suited objectives to learn time-dependent quantile functions to optimize the entire path distribution, or designing more sophisticated conditional quantile functions for tasks like class-conditional or text-to-image generation.
Acknowledgments.
Appendix A Examples of One-Dimensional Flows
We provide three interesting examples, namely the well-established diffusion flow, the recently proposed Kac flow, and the Wasserstein gradient flow of the MMD functional with negative absolute distance kernel towards a uniform measure. Paths of the processes are depicted in Figure 7 and their probability flows are shown in Figures 8, 9 and 10.
In each case, the absolutely continuous curve $\mu_{t}$ starting in $\delta_{0}$ (e.g. conditional) and the corresponding velocity field can be given analytically. Note that in the latter two cases, multi-dimensional generalizations of the flows are not trivially given, which further underlines the strength of our 1D approach. Henceforth, if the measures $\mu_{t}$ admit a density function, we will denote it by $p_{t}$.